BRICS Notes Series, 2002

July 7, 2003

This document is also available as PostScript and DVI.

References

[NS-02-8] Abstract, PostScript, PDF.

Peter D. Mosses, editor. Proceedings of the Fourth International Workshop on Action Semantics, AS 2002, (Copenhagen, Denmark, July 21, 2002), December 2002. vi+133 pp.

[NS-02-7] Abstract, PostScript, PDF.

Anders Mller. Document Structure Description 2.0. December 2002. 29 pp.

[NS-02-6] Abstract, PostScript, PDF, DVI.

Aske Simon Christensen and Anders Mller. JWIG User Manual. October 2002. 35 pp.

[NS-02-5] Abstract, PDF.

Patrick Cousot, Lisbeth Fajstrup, Eric Goubault, Maurice Herlihy, Martin Rauen, and Vladimiro Sassone, editors. Preliminary Proceedings of the Workshop on Geometry and Topology in Concurrency Theory, GETCO ’02, (Toulouse, France, October 30–31, 2002), October 2002. vi+97.

[NS-02-4] Abstract, PostScript, PDF.

Daniel Gudbjartsson, Anna Inglfsdttir, and Augustin Kong. An BDD-Based Implementation of the Allegro Software. August 2002. 2 pp.

[NS-02-3] Abstract, PDF.

Walter Vogler and Kim G. Larsen, editors. Preliminary Proceedings of the 3rd International Workshop on Models for Time-Critical Systems, MTCS ’02, (Brno, Czech Republic, August 24, 2002), August 2002. vi+141 pp.

[NS-02-2] Abstract, PDF.

Zoltn sik and Anna Inglfsdttir, editors. Preliminary Proceedings of the Workshop on Fixed Points in Computer Science, FICS ’02, (Copenhagen, Denmark, July 20 and 21, 2002), June 2002. iv+81 pp.

[NS-02-1] Abstract, PDF. Anders Mller and Michael I. Schwartzbach. Interactive Web Services with Java: JSP, Servlets, and JWIG. April 2002. 99 pp.